# Direct observation of the crystal electric-field splitting under magnetic field and uncovering field-induced magnetic phase transition in triangular rare-earth magnet CsErSe2

### v1

@@ Abstract
An indispensable step toward understanding magnetic interaction in rare-earth magnets is the determination of spatially anisotropic single-ion properties resulting from the crystal electric field (CEF) physics. The CEF Hamiltonian exhibits a discrete energy spectrum governed by a set of independent parameters that reflect the site symmetry of the magnetic ion. However, experimentally determining these parameters for magnetic ions at low-symmetry sites has been proven highly challenging. In this study, we directly measured the CEF level splitting under magnetic fields ($\mathbf{B}$) using optical spectroscopy and extracted both CEF parameters and the exchange energies of a triangular insulating magnet CsErSe2 as a model system. With increasing field, we find many CEF levels undergo level-crossing which accompanies switching of the eigenstate. Particularly, such a crossing occurring at the ground state results in a step-like increase in magnetization that we captured with the low-temperature AC magnetic susceptibility measurements. Our work demonstrates that the accurately determined CEF Hamiltonian parameters enable uncovering the rich physics of field-induced collective magnetic phenomena, and potentially lead to a new route to the emergence of magnetic frustration.
@@ I Introduction
Magnetic insulators with strong spin-orbit coupling are now widely recognized as a robust platform to investigate a range of quantum many-body phases, including quantum spin liquids (QSLs) and other phenomena, via various forms of magnetic frustration Savary and Balents (2016); Scheie et al. (2019).
In particular, compounds based on $4f$ rare-earth (RE) ions offer an excellent playground to explore various interesting magnetic interactions, due to their unique combination of geometric frustration arising from the two-dimensional triangular lattice and multiple magnetic degrees of freedom induced by spin-orbit coupling (SOC) and crystal electric field (CEF) interaction. Gardner et al. (2010); Rau and Gingras (2019); Bordelon et al. (2019); Scheie et al. (2024); Paddison et al. (2017); Gao et al. (2018); Li et al. (2015); Mauws et al. (2018); Scheie et al. (2020).
In RE-based insulating magnets, however, detailed understanding of the microscopic physics is a prerequisite for developing effective spin models that serve as a basis for theories of many-body phenomena. For a free RE ion, the spin ($S$) and orbital ($L$) angular momenta of unpaired electron couple into a total angular momentum $J=L+S$, which usually turns out to be a good quantum number. When a RE ion is placed in a crystalline environment, the Coulomb interaction with surrounding ligands generates an electrostatic crystal electric field (CEF), breaking the inherent ($2J+1$)-fold degeneracy. Stevens first developed a theoretical approach to describe this phenomenon in 1952 Stevens (1952), introducing angular momentum operators known as Stevens operators to express the ion’s electrostatic potential and directly obtaining the $g$ tensor. This work was later expanded further by Hutchings, who calculated crystal field levels using a simplified point-charge ionic model Hutchings (1964). Consequent studies established that the main characteristic of the CEF Hamiltonian parameters (i.e. Stevens’ operators and their coefficients) is determined solely by the site symmetry of a single RE ion – for example, two or four parameters for cubic, six or eight for hexagonal, and up to twenty-six for lower symmetry environments.
@@
In this work, we directly measure the magnetic field $(\mathbf{B})$ dependence of the Zeeman splittings of the CEF levels of CsErSe2 using magneto-topical spectroscopy. From these results, we determine the parameters for the Hamiltonian, i.e. Stevens’ coefficients and the exchange energy scales, to obtain the full field-dependent CEF spectra in two high symmetry directions parallel to the crystalline axes $\hat{b}$ and $\hat{c}$, as shown in Fig. 1 (a) and (b). We also predict the field-induced transition that originates from the crossing of the ground state in $\mathbf{B}\parallel c$ geometry, at $B_{C}\approx 5.4$ T, where magnetization along the $c$-axis displays the field-dependence similar to a metamagnetic transition. Fig. 1(e) shows the calculated magnetization. We verified this transition with the field dependence of AC magnetic susceptibility down to $T=25$ mK, showing that this feature remains robust even below $T_{N}=110$ mK, where long-range ordering sets in at zero field.
@@
Figure 1: CEF energy spectra of single Er3+  ion in CsErSe2 under applied magnetic field (a) $\mathbf{B}\parallel b$ and (b) $\mathbf{B}\parallel c$, calculated from Eq. (1) using the obtained parameters from this study. Eight doublets at zero field ($B=0$) (shaded area on the left side of each plot) are clustered such that five lower levels lie below around 5 meV and the rest above 23 meV. Note that the ground state wave function changes due to the level crossing at $B_{C}=5.4$ T, and another crossing between the upper Zeeman-split level of the ground state and the lower of the first excited occurs at $B=1.4$ T in $\mathbf{B}\parallel c$ shown in (b) (c) Crystal structure of CsErSe2 adopts $P6_{3}/mmc$ space group Xing et al. (2020) (d) Er3+(red) triangular layers formed by edge-sharing ErSe6 octahedra, exhibiting $D_{3d}$ site symmetry (e) The calculated magnetization $M_{b}$ ($\mathbf{B}\parallel b$) and $M_{c}$ ($\mathbf{B}\parallel c$) with mean-field approximation (see the text) at $T=$ 25 mK are shown a function of applied field, which will be discussed in Sec.III.1 and Sec.III.2. The inset displays the same quantities at lower field range with $M_{b}$ and $M_{c}$ calculated at $T=2$ K added. Magnetization curves calculated with the MF approximation at different temperatures are Fig. 3 in the Appendix.
Our work reveals the rich physics of CEF interactions under high magnetic fields, demonstrating that even single-ion physics can host unexpected complexity. This opens a new avenue for exploring the frustrated magnetism unique to rare-earth magnets, where the CEF energy scales are small enough to be manipulated with moderate magnetic field intensity available in a modern laboratory.
@@
The paper is organized as follows. Sec.II describes the CsErSe2 samples, experimental methods, and the Hamiltonian used to calculate energy spectra and magnetic properties. Sec.III.1 presents magneto-optical spectroscopy and analysis for two field directions, from which we determine the Stevens coefficients and spin-exchange energies. Sec.III.2 reports AC susceptibility and magnetization measurements compared with calculations, while Sec.III.3 examines the long-range magnetic order revealed by elastic neutron scattering. Sec.IV discusses the comparison with calculated intensities and the overall implications of our work, and Sec.V summarizes the study. Additional data and analyses are provided in the Appendices.
@@ II.1 Materials and Experimental Methods
CsErSe2 crystallizes in a $P6_{3}/mmc$ space group Xing et al. (2020); Scheie et al. (2020) as illustrated in Fig. 1(c). The triangular-lattice planes [Fig. 1(d)] consist of edge-sharing ErSe6 octahedra, allowing $D_{3d}$ site symmetry for Er3+. Within the $ab$-plane, the Se atoms mediate antiferromagnetic (AFM) superexchange interactions between nearest-neighbor Er3+ion pairs, resulting in frustrated triangular magnetism. We identify CsErSe2 indeed exhibits long-range order at zero field at $T_{N}=110$ mK, which we discuss in details in Sec.III.3.
@@
Magneto-infrared spectroscopy was performed in a 17.5 T superconducting magnet with a Fourier-transform infrared (FTIR) spectrometer by Bruker Vertex 80v. Broadband IR light provided by the FTIR spectrometer is guided via evacuated brass light pipes to the sample space, situated at the center of the magnet. Transmitted IR light is delivered to a composite Si-bolometer (IR Labs, Inc.) mounted in the same cryogenic environment, cooled by low-pressure helium gas to a temperature of approximately 5 K. We used both Faraday ($\mathbf{B}\parallel c$) and Voigt ($\mathbf{B}\parallel b$) geometries with the wave vector of the incoming IR light to measure the CsErSe2 transmission spectra. We find that CsErSe2 completely absorbs the spectral range of $130\leq E\leq 210$ cm-1.
@@
We measured the neutron diffraction spectra at the HB2A spectrometer at ORNL’s High Flux Isotope Reactor (HFIR) Garlea et al. (2010). We measured 3 g of loose powder of CsErSe2 in a copper can mounted in a dilution refrigerator at temperatures $T=0.5$ K and $T=0.05$ K using $\lambda=2.41$ Å  neutrons. The data are plotted in Fig. 4. The magnetic scattering was isolated by subtracting the high-temperature data from the low-temperature data. We performed Reitveld refinements to the neutron data using the Fullprof software package Rodríguez-Carvajal (1993).
@@ II.2 Single ion Hamiltonian
In CsErSe2 , the first term can be expressed as a linear combination of the six symmetry-allowed Stevens operators, $\hat{O}^{m}_{n}$, for the $D_{3d}$ site symmetry of the ErSe6 octahedral environment Scheie et al. (2020); Ding et al. (2023)
@@
At zero field (ZF), the energy eigenstate of this Hamiltonian splits the $J=15/2$ multiplet into eight Kramers doublets, $|n_{\pm}\rangle$ with $n=0,1,2,\cdots,7,8$ [shaded area of Fig. 1(a) and (b)]. Upon applying magnetic field, each doublet splits into two states according to the Zeeman interaction:
@@
where the Landé $g$-factor, $g_{J}=6/5$, is used for Er3+moments.
@@
Figure 2: (a,b) Field dependence of normalized FIR absorption intensity at $T=5$ K below 120 cm-1and magnetic field applied in $\mathbf{B}\parallel b$ (a) and $\mathbf{B}\parallel c$ (b). (c) Normalized peak intensity of Raman shift, denoted as $E$, as a function of field in $\mathbf{B}\parallel c$ at $T=5$ K. The FIR spectra in (a)and (b) do not show any field-independent features due to subtracting the ZF spectra. The Raman spectrum, however, displays the field-independent Raman active phonon mode at 49.3 cm-1, which corresponds to the $E_{2g}^{1}$ mode (See Table 2 in Appendix ). The peak positions in (a), (b) and (c) are marked with circles in (d), triangles (e) and crosses in (f). Solid lines in (a-f) indicate transitions from the ground state to field-split excited CEF levels. 1(a-b). Gray-shaded areas mark the regions $B_{\rm ext}>0$ and $E>0$ for clarity.
@@
This is determined self-consistently via $B=B_{\rm ext}+\lambda M(B)$, where $B_{\rm ext}$ refers to externally applied field and we define the MF parameter $\lambda=\frac{q\mathcal{J}}{g_{J}\mu_{B}}$, where $\mathcal{J}=\mathcal{J}_{\perp}$ or $\mathcal{J}_{z}$, depending on the applied magnetic field orientations Blundell (2001).
@@ III.1 Magneto-optical spectroscopy
Leveraging the field-dependent Zeeman splittings, we demonstrate a high-fidelity determination of the CEF parameters in CsErSe2 . Fig. 2 displays our main result of magneto-optical spectroscopy, shown below 120 cm-1($\approx$ 15 meV). FIR absorption up to 17.5 T are shown in (a) and (b) for the configurations of $\mathbf{B}\parallel b$ and $\mathbf{B}\parallel c$, respectively. Panel (c) shows the Raman shift under a magnetic field up to 14 T in $\mathbf{B}\parallel c$ geometry. Both IR and Raman data were obtained at the bath temperature at $T=5$ K. For clarity, FIR spectra in (a) and (b) are shown with the ZF spectrum subtracted and normalized to the global maximum, to show only field-dependent CEF levels, while the Raman shift as a function of field, shown in panel (c) is normalized only by the global maximum intensity.
@@
FIR spectra of Fig.2 (a) and (b) display a myriad of crisscrossing energy levels with distinct field dependencies, highly dependent on field orientation, highlighting the strong magneto-anisotropy.
@@
Despite fundamentally different natures of FIR and Raman spectra, the resemblance shown in panels (b) and (c), both taken in $\mathbf{B}\parallel c$, is remarkable, clearly capturing the same physics, except for the field-independent peaks at $E\approx 49$ cm-1in panel (c). It is identified as $E_{2g}^{1}$ phonon from the phonon band calculation (See Table 2 in Appendix).
@@
As mentioned in Sec.II, the FIR spectra of CsErSe2  suffer from complete absorption in $125\leq E\leq 210$ cm-1, above which few discernible field-dependent features were observed. Raman data, however, display discernible features up to $500$ cm-1[ See Fig. 6(a) in Appendix], where transition intensities involving the three highest energy doublets [Fig. 1 (b)] do not appear. This indicates the corresponding intensity is far weaker than those below $E<120$ cm-1, due to exceedingly smaller transition rates. We will return this in Sec. IV.
@@
To obtain the seven parameters at a given field-orientation, we consider only the five lowest doublets, as shown in Fig. 2(d-f). We fit the Raman spectra in panel (f) first with seven parameters – six Stevens’ operators coefficients and $\mathcal{J}_{z}$, and then fit the FIR spectra in $\mathbf{B}\parallel b$ with fixed Stevens Operators’ coefficients to find $\mathcal{J}_{\perp}$. Table 1 lists parameters obtained from the fitting, which agree remarkably well with previous inelastic neutron scattering results Scheie et al. (2020). This is unique to Er3+, a Kramers ion with high $J$ to have seven gaps at ZF that enhance the precision of the fitting Scheie (2021). Using these parameters, we calculated the field evolution of the CEF levels for both field orientations as seen in Fig. 1 (a) and (b). The solid lines shown in Fig. 2 (a-f) are computed by simply subtracting the ground state energy from each excited state energy. This simple procedure alone already captures the peak locations remarkably well, confirming that most features in our spectroscopy data originate from the splitting of the CEF levels.
@@
$\mathcal{J}_{z}$ $2.64\times 10^{-3}$ $2.4\times 10^{-2}$
$\mathcal{J}_{\perp}$ $0.53\times 10^{-3}$ $0.2\times 10^{-2}$
We now focus on the $\mathbf{B}\parallel c$ spectra, comparing the calculated energy spectra in Fig. 1 (b)] with our data. The most striking feature revealed by calculation is the level crossing on the ground state $B_{C}\approx 5.4$ T, and the crossing of the first excited state at $B_{C}*\approx 1.4$ T. These features are too low in energy for any spectroscopic probe and were not observed in Fig.2(b-c). These level crossings are expected due to the narrowly spaced CEF levels of Er3+even at ZF. For example, at zero field, the gap between ground and the lowest excited state doublets is found to be 0.67 meV, and as the field in $\mathbf{B}\parallel c$, the non-monotonic energy difference between the ground and the first excited state do not exceed this value until the field increases above 12 T.
The direct consequence of the crossing in the ground state is an abrupt change in magnetization, calculated with the MF approximation, as shown in Fig. 1(e): The magnetization along the $c$-axis ($M_{c}$) exhibits a step-like rise at around $B_{C}$, resembling a metamagnetic transition Jacobs and Lawrence (1967). Meanwhile, $M_{b}$, the magnetization along the $b$-axis does not have such a feature. With increasing temperature, the step-like rise in $M_{c}$ disappears quickly, due to increasing thermal populations according to the Boltzmaan factor of the excited states, one of characteristic of the CEF levels.

### v2

# Uncovering field-induced magnetic phase transition by direct observation of the crystal electric-field splitting in a rare-earth magnetic insulator

@@ Abstract
An indispensable step toward understanding magnetic interactions in rare-earth magnets is to determine the spatially anisotropic single-ion properties set by the crystal electric field (CEF) physics. The CEF Hamiltonian yields a discrete energy spectrum governed by a set of parameters reflecting the local site symmetry of the magnetic ion. However, experimentally determining these parameters, especially for ones at low-symmetry sites remains highly challenging. In this work, we directly measure the CEF level splitting of CsErSe2 under magnetic fields using optical spectroscopy. This enables us to determine the CEF parameters and to predict the metamagnetic-like transition arising from a level crossing at the ground state. We also identify another crossing in the first excited state that leads to a non-monotonic energy gap and hence, strongly influences the temperature and field dependence of the magnetic properties. Our results highlight the capacity of single-ion physics to drive rich and unanticipated phenomena in rare-earth magnetic insulators under applied magnetic field.
@@ I Introduction
Magnetic insulators with strong spin-orbit coupling are now widely recognized as a robust platform to investigate a range of quantum many-body phases, including quantum spin liquids (QSLs) and other phenomena, via various forms of magnetic frustration Savary and Balents (2016); J. et al. (2019).
In particular, compounds based on $4f$ rare-earth (RE) ions provide an excellent platform for exploring a wide range of magnetic interactions. These systems combine geometric frustration from the two-dimensional triangular lattice with multiple magnetic degrees of freedom originating from strong spin-orbit coupling (SOC) and crystal electric field (CEF) effects. Gardner et al. (2010); Rau and Gingras (2019); Bordelon et al. (2019); Scheie et al. (2024); Paddison et al. (2017); Gao et al. (2018); Li et al. (2015); Mauws et al. (2018); Scheie et al. (2020).
In RE-based insulating magnets, however, a detailed understanding of the microscopic physics is a prerequisite for developing effective spin models that serve as a basis for theories of many-body phenomena. For a free RE ion, the spin ($S$) and orbital ($L$) angular momenta of unpaired electrons couple into a total angular momentum $J=L+S$, which becomes a good quantum number. When a RE ion is placed in a crystalline environment, the Coulomb interaction with surrounding ligands generates an electrostatic crystal electric field (CEF), lifting the ($2J+1$)-fold degeneracy. For half-odd-integer $J$, the result is a set of Kramers doublets. At sufficiently low temperatures, a restriction of the dynamics to the lowest such doublet may be invoked to justify using an effective pseudospin-1/2 model Gao et al. (2018); Sibille et al. (2015); Bordelon et al. (2019); Dai et al. (2021). From an experimental standpoint, this allows a significant diversification of both materials and magnetic phe- nomena, while on the theoretical side a minimal model may be relatively simple, or even a realization of one of the paradigm models in frustrated magnetism.
Stevens first developed a theoretical framework for describing crystal-field effects in 1952 Stevens (1952), introducing the operator-equivalent formalism in which the CEF Hamiltonian with a given $J$ multiplet is written as $\sum B_{m}^{n}\hat{O}_{m}^{n}(\mathbf{J})$, where $\hat{O}_{m}^{n}$’s are known as Stevens’ operators and $B_{m}^{n}$’s are the corresponding coefficients. Hutchings later extended this approach by calculating split energy levels using a simplified point-charge model Hutchings (1964). Subsequent studies established that the allowed Stevens operators ($O_{m}^{n}$’s) and their coefficients ($B_{m}^{n}$’s) in the CEF Hamiltonian are determined entirely by the local site symmetry of the rare-earth ion. When combined with the Zeeman interaction, this formulation allows direct calculation of the anisotropic $g$-tensor of the ground states once the CEF parameters ($B_{m}^{n}$’s) are known.
@@
In this work, we directly measure the magnetic field $(\mathbf{B})$ dependence of the Zeeman splittings of the CEF levels of CsErSe2 using far-infrared and Raman magneto-optical spectroscopy. From these results, we determine the parameters for the Hamiltonian, i.e. Stevens’ coefficients and the exchange energy scales, to obtain the full field-dependent CEF spectra in two high symmetry directions along the crystalline axes $\hat{b}$ and $\hat{c}$, as shown in Fig. 1 (a) and (b). Using these results, we predict the field-induced magnetic transition that arises from the crossing of the ground state in $\mathbf{B}\parallel c$ geometry, at $B_{C}\approx 5$ T, where magnetization along the $c$-axis displays the field-dependence similar to a metamagnetic transition. Fig. 1(e) shows the calculated magnetization. We verified this transition with the field dependence of AC magnetic susceptibility down to $T=25$ mK, showing that this feature remains robust even below $T_{N}=110$ mK, where long-range ordering sets in at zero field.
@@
Figure 1: The field dependent CEF energy levels of single Er3+  ion in CsErSe2 under applied magnetic field (a) $\mathbf{B}\parallel b$ and (b) $\mathbf{B}\parallel c$, calculated from Eq. (1) using the obtained parameters from this study. Eight doublets at zero field ($B=0$) (shaded area on the left side of each plot) are clustered such that five lower levels lie below around 5 meV and the rest above 23 meV. Note that the ground state wave function changes due to the level crossing at $B_{C}\approx 5$ T, and another crossing between the upper Zeeman-split level of the ground state and the lower of the first excited occurs at $B_{C}^{*}\approx 1.4$ T in $\mathbf{B}\parallel c$ shown in (b) (c) Crystal structure of CsErSe2 adopts $P6_{3}/mmc$ space group Xing et al. (2020) (d) Er3+(red) triangular layers formed by edge-sharing ErSe6 octahedra, exhibiting $D_{3d}$ site symmetry (e) The calculated magnetization $M_{b}$ ($\mathbf{B}\parallel b$) and $M_{c}$ ($\mathbf{B}\parallel c$) with mean-field approximation (see the text) at $T=$ 25 mK are shown a function of applied field, which will be discussed in Sec. III.1 and Sec. III.2. The inset displays the same quantities at lower field range with $M_{b}$ and $M_{c}$ calculated at $T=2$ K added. Magnetization curves calculated with the MF approximation at different temperatures are Fig. 3 in the Appendix.
Our work reveals unexpectedly complex magnetic properties of CsErSe2 , driven by single-ion physics arising from multiple CEF level crossings under applied magnetic field. This opens a new avenue for exploring magnetic properties unique to the $4f$ rare-earth insulating magnets, where the smaller CEF energy scales, compared to their $3d$-counterpart, can be readily tuned with moderate intensity of magnetic field.
@@
The paper is organized as follows. Sec. II describes the CsErSe2 samples, experimental methods, and the Hamiltonian used to calculate energy spectra and magnetic properties. Sec. III.1 presents magneto-optical spectroscopic data obtained from far-infrared and Raman measurements that were used to determine the Stevens coefficients and spin-exchange energies. Sec. III.2 reports AC susceptibility and magnetization measurements compared with calculations, while Sec. III.3 examines the long-range magnetic order revealed by elastic neutron scattering. Sec. IV discusses the comparison with calculated intensities and the overall implications of our work, and Sec. V summarizes the study. Additional data and analyses are provided in the Appendices.
@@ II.1 Materials and Experimental Methods
CsErSe2 crystallizes in the $P6_{3}/mmc$ (#194) space group Xing et al. (2020); Scheie et al. (2020), shown in Fig. 1(c). The triangular-lattice planes [Fig. 1(d)] consist of edge-sharing ErSe6 octahedra, allowing $D_{3d}$ site symmetry for Er3+. Within the $ab$-plane, the Se atoms mediate antiferromagnetic (AFM) superexchange interactions between nearest-neighbor Er3+ion pairs. We find CsErSe2 indeed exhibits long-range order at zero field at $T_{N}=110$ mK, which we discuss in detail in Sec. III.3.
@@
Far-infrared magneto-spectroscopy was performed using a Fourier-transform infrared spectrometer (Bruker Vertex 80v) coupled to a 17.5 T vertical-bore superconducting magnet. Broadband IR light provided by the mercury lamp is guided via evacuated brass light pipes to the sample space, situated at the center of the magnet. The intensity of the radiation transmitted through the sample was detected by a composite Si bolometer (IR Labs, Inc.) mounted in the same cryogenic environment and cooled by low-pressure helium gas to approximately 5.5 K. The FIR transmission spectra of CsErSe2 were measured for the Faraday geometry ($\mathbf{B}\parallel c$) and Voigt geometry ($\mathbf{B}\parallel b$). To highlight the field-dependent absorption modes, the spectrum at each magnetic field was divided by a reference spectrum defined as the average over all measured fields. The resulting normalized transmission is sensitive only to intensity changes induced by the magnetic field and eliminates contributions from field-independent vibrational absorption and instrumental artifacts. No detectable transmission in the range of 130210130-210 cm-1 due to strong absorption in CsErSe2 .
@@
We measured the neutron diffraction spectra at the HB2A spectrometer at ORNL’s High Flux Isotope Reactor Garlea et al. (2010). We measured 3 g of loose powder of CsErSe2 in a copper can mounted in a dilution refrigerator at temperatures $T=0.5$ K and $T=0.05$ K using $\lambda=2.41$ Å  neutrons. The data are plotted in Fig. 4. The magnetic scattering was isolated by subtracting the high-temperature data from the low-temperature data. We performed Reitveld refinements to the neutron data using the Fullprof software package Rodríguez-Carvajal (1993).
Figure 2: (a,b) Field dependence of normalized FIR absorption spextra at $T=5.5$ K below 120 cm-1and magnetic field applied in $\mathbf{B}\parallel b$ (a) and $\mathbf{B}\parallel c$ (b). (c) Raman shift, denoted as $E$, as a function of field in $\mathbf{B}\parallel c$ measured at $T\approx 5$ K are shown. The field-independent Raman active phonon mode at 49.3 cm-1corresponds to the $E_{2g}^{1}$ mode is clearly visible (See Table 2). The peak positions in (a), (b) and (c) are marked with circles in (d), triangles (e) and crosses in (f). Solid lines in (a-f) indicate transitions from the ground state to field-split excited CEF levels. 1(a-b). Gray-shaded areas mark the regions $B_{\rm ext}>0$ and $E>0$ for clarity.
@@ II.2 Single ion Hamiltonian
In CsErSe2 , the first term can be expressed as a linear combination of the six symmetry-allowed Stevens operators, $\hat{O}_{m}^{n}$, for the $D_{3d}$ site symmetry of the ErSe6 octahedral environment Scheie et al. (2020); Ding et al. (2023)
@@
At zero field (ZF), the energy eigenstate of this Hamiltonian splits the $J=15/2$ multiplet into eight Kramers doublets, $|n_{\pm}\rangle$ with $n=0,1,2,\cdots,7$ [shaded area of Fig. 1(a) and (b)]. Upon applying magnetic field, each doublet splits into two states according to the Zeeman interaction:
@@
where the Landé $g$-factor, $g_{J}=6/5$, is used for Er3+ ion.
@@
This is determined self-consistently via $B=B_{\rm ext}+\lambda M(B)$, where $B_{\rm ext}$ refers to externally applied field and we define the MF parameter $\lambda=\frac{q\mathcal{J}}{g_{J}\mu_{B}}$, where $g_{J}$ refers Landé $g$-factor, $\mathcal{J}=\mathcal{J}_{\perp}$ or $\mathcal{J}_{z}$, depending on the applied magnetic field orientations Blundell (2001).
@@ III.1 Magneto-optical spectroscopy
The great advantage of magneto-optical spectroscopy is its ability to determine the CEF parameters with high fidelity by leveraging the Zeeman splitting of Kramers doublets. Fig. 2 displays our main result of magneto-optical spectroscopy, shown below 120 cm-1($\approx$ 15 meV). The heat map of normalized FIR absorption spectra up to 17.5 T are shown in (a) and (b) for the configurations of $\mathbf{B}\parallel b$ and $\mathbf{B}\parallel c$, respectively. Panel (c) shows the Raman shift under a magnetic field up to 14 T in $\mathbf{B}\parallel c$ geometry. Both FIR and Raman data were obtained at the bath temperature at $T=5.t$ K. Both FIR and Raman spectra shown are normalized with global maximum intensity so that they are displayed in the sam color scale.
@@
FIR spectra in Fig. 2(a) and (b) display a myriad of crisscrossing with distinct field dependencies that are highly sensitive to field orientation, highlighting the strong magneto-anisotropy.
@@
Despite fundamentally different natures of FIR and Raman spectra, the resemblance shown in panels (b) and (c), both taken in $\mathbf{B}\parallel c$, is remarkable, clearly capturing the same excitations between the CEF energy levels, except for the field-independent peaks at $E\approx 49$ cm-1in panel (c). It is identified as $E_{2g}^{1}$ phonon from the phonon band calculation (See Table 2 in Appendix).
@@
The FIR spectra of CsErSe2  suffer from complete absorption in approximately 127210127-210 cm-1, and above 210 cm-1, few discernible field-dependent features were observed. Raman data, however, display three field-dependent features up to $500$ cm-1[Fig. 6(a) in Appendix], but they are far weaker in intensity, indicating the three highest energy doublets exhibit exceeldingly smaller transition rate. We will return this in Sec. IV.
@@
To obtain the seven parameters at a given field-orientation, we consider only the five lowest doublets, as shown in Fig. 2(d-f). We fit the Raman spectra in panel (f) first with seven parameters – six Stevens’ operators coefficients and $\mathcal{J}_{z}$, and then fit the FIR spectra in $\mathbf{B}\parallel b$ with fixed Stevens operators’ coefficients to find $\mathcal{J}_{\perp}$ Scheie (2021). Table 1 lists parameters obtained from the fitting, which agree well with previous inelastic neutron scattering results Scheie et al. (2020). Obtaining the CEF parameters with high precision with ZF data could be unique to Er3+, due to its multiple CEF levels within a narrow energy range. Using these parameters, we calculated the field evolution of the CEF levels for both field orientations as seen in Fig. 1 (a) and (b). The solid lines shown in all panels of Fig. 2 are computed by simply subtracting the ground state energy from each excited state energy. This simple procedure alone already captures the peak locations remarkably well, confirming that most features in our spectroscopy data originate from electronic transitions from the ground to the higher-energy CEF levels.
@@
$\mathcal{J}_{z}$ $2.64\times 10^{-3}$ $2.4\times 10^{-3}$
$\mathcal{J}_{\perp}$ $0.53\times 10^{-3}$ $0.2\times 10^{-3}$
We now focus on the CEF transitions for $\mathbf{B}\parallel c$, comparing the calculated energy spectra in Fig. 1 (b)] with our data. The most striking feature revealed by calculation is the level crossing on the ground state $B_{C}\approx 5$ T, and the crossing of the first excited state at $B_{C}^{*}\approx 1.4$ T. These features are too low in energy for any spectroscopic probe and were not observed in Fig.2(b-c). These level crossings are expected due to the narrowly spaced two lowest Kramers doublets of Er3+, 0.67 meV apart at ZF. Upon applying field in $\mathbf{B}\parallel c$, the non-monotonic field dependent energy gap between the ground and the first excited state emerges [Fig. 3(a)], which dictates the field and temperature dependences of magnetizations.
The direct consequence of the crossing in the ground state is an abrupt change in magnetization, calculated with the MF approximation, as shown in Fig. 1(e): The magnetization along the $c$-axis ($M_{c}$) exhibits a step-like rise at around $B_{C}$, resembling a metamagnetic transition Jacobs and Lawrence (1967). Meanwhile, $M_{b}$, the magnetization along the $b$-axis does not have such a feature. With increasing temperature, the step-like rise in $M_{c}$ becomes smeared out due to Boltzmann thermal population of the low-lying excited CEF states, which will be discussed in the next Sec. III.2.